stefani bought enough powered juice mix to make eight and three fourths gallons of juice if her kids go through one and one fourths gallons of juice per week how many weeks of juice did she buy?
Answer by checkley77(12844)

(8 3/4)/(1 1/4)
(35/4)/(5/4)
35/4*4/5=35/5=7 weeks.
Proof:
8.75/1.25=7